OpenStudy (anonymous):

Find the value of the constant c that makes the following function continuous on (−∞,∞). f(y)={cy+8cy2−8if y≤5if y>5 c=

OpenStudy (anonymous):

\(\LaTeX\) time

OpenStudy (dumbcow):

wait what is function when y>5 ??

OpenStudy (anonymous):

\[f(y) = \left\{\begin{array}{rcc} cy+8 & \text{if} & y\leq 5 \\ cy^2-8& \text{if} & y>5 \end{array} \right. \]

OpenStudy (dumbcow):

oh didn't catch that ... anyway plug in y=5 and set both cases equal

OpenStudy (anonymous):

i tryed but it saey incorrect

OpenStudy (dumbcow):

5c +8 = 25c -8 this is what you did? what did you get for c?

OpenStudy (anonymous):

thinks i got it

OpenStudy (anonymous):

4/5

OpenStudy (dumbcow):

correct :)